OpenStudy (anonymous):

Describe the transformation of the graph of f into the graph of g as either a horizontal or vertical stretch. f as a function of x is equal to the square root of x and g as a function of x is equal to the square root of zero point five x

OpenStudy (mathmate):

f(x)=sqrt(x) g(x)=sqrt(0.5x) is a horizontal with a factor of 2 = sqrt(0.5) sqrt(x) is a vertical stretch with a factor of sqrt(0.5)=sqrt(2)/2 Since stretching3